OpenStudy (anonymous):

HELP PLEASE. For what values of x and y are the triangles congruent to HL? Drawing below.

OpenStudy (anonymous):

What's HL?

OpenStudy (anonymous):

hmmm.. that makes it hard to answer

OpenStudy (anonymous):

Yeah, but I'm trying to figure out what HL is..

OpenStudy (anonymous):

I hate question's like these that aren't very specific

OpenStudy (anonymous):

Its talking about the HL theorem

OpenStudy (anonymous):

H I think hypotenuse and L is length

OpenStudy (anonymous):

Oh yeah!

OpenStudy (anonymous):

then set them equal and solve

OpenStudy (anonymous):

Could you show me how please?

OpenStudy (anonymous):

sorry, I looked at it, let me think on this! Sorry, i didn't see it had 2 variables

OpenStudy (anonymous):

Oh okay!

OpenStudy (anonymous):

What math is this?

OpenStudy (anonymous):

take longer to solve I will get back to you later

OpenStudy (anonymous):

oh I did geom last year, let me see if I remember this

OpenStudy (anonymous):

Ok thanks!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Ans. x = -1, y=3 Solution: You basically need to prove triangles congruent by HL i.e. by Hypotenuse & Length(length should probably represent height) So in the first triangle, Hypotenuse = 3y +x & Length = y-x in 2nd triangle, H = y+5 & L = x+5 For concurrency by HL, both the H's and L's shud be equal respectively i.e. 3y+x = y+5 And y-x = x+5 By solving these eqns u will get, x=-1 and y=3 Trivia: Hypotenuse is basically the longest side of the triangle and is opposite to the right angle.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

sure, i am right now on call :P wait-up!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

for H, 3y+x=y+5 2y+x=5 ~ equation(1) for L, y-x=x+5 y-2x=5 y=5+2x ~ equation(2) put this value of y in eqn 1 u get, 2(5+2x)+x = 5 10+4x+x=5 5x=5-10 5x=-5 x=-1 now put this in eqn 2 u get, y=5+2(-1)=5-2=3 Ans. x=-1 & y=3

OpenStudy (anonymous):

Oh wait so this is for the first triangle?

OpenStudy (anonymous):

this is for both the triangles. For these triangles to be congruent x should be = -1 and y should be = 3

OpenStudy (anonymous):

Oh ok thanks!!

OpenStudy (anonymous):

Hypotenuse Leg